The electric force experienced by a charge of 1.0 × 10-8 C is 1.5 × 10-3 N. Find the magnitude of the electric field at the position of the charge.

Grade:upto college level

1 Answers

Navjyot Kalra
9 years ago
Sol. F base e = 1.5 × 10^–3 N, q = 1 × 10^–6 C, F base e = q × E ⇒ E = F base e/q = 1.5 * 10^-3/1 * 10^-6 = 1.5 * 10^3 N/C